Research objective: The population mean is not
equal to 1,500.
σ = 220, n = 125, x = 1525
1
Expert's answer
2020-12-22T19:40:13-0500

"Z_{value}= \frac {\mu_X-1500} {\sigma_X}= \frac {1525 - 1500} {220/ \sqrt{125}}= 1.27\\
Z_{0.025} = -1.96\\
Z_{0.975} = 1.96\\
Z_{value} \in (-1.96,1.96)"


Hence, we cannot reject the null hypothesis that the population mean is equal to 1,500 at significant level of 0.05.